An object is launched at a velocity of 20 m/s in a direction making an angle of 25° upward with the horizontal
1
Expert's answer
2020-04-21T19:03:06-0400
"h=\frac{(v\sin{25})^2}{2g}=\frac{(20\sin{25})^2}{2(9.8)}=3.64\ m"

"t=\frac{2(v\sin{25})}{g}=\frac{2(20\sin{25})}{(9.8)}=1.72\ s"

"R=v\cos{25}t=20\cos{25}(1.72)=31.26\ m"